# Neural Bridge Processes

### v1

@@ 1. The ELBO decomposes into:
2. Sample from $q(y_{t}|y_{0},x)$: Using the forward process properties by Equation (10) in the main text, we can write:
@@ 5. Final Noise-Prediction Loss: The KL terms simplify to a weighted $L_{2}$ loss on the noise:
This proves Equation (18) in the main text, and the denoising network $\epsilon_{\theta}(y_{t},t,x)$ is self-consistent with respect to the condition on $x$.
@@ Appendix C Background: Neural Processes
Neural Processes (NPs) Garnelo et al. (2018b; a); Kim et al. (2019); Louizos et al. (2019) combine the expressiveness of neural networks with the probabilistic reasoning of Gaussian Processes (GPs) Rasmussen (2003). While GPs offer principled uncertainty quantification, they suffer from poor scalability Snelson & Ghahramani (2005); Titsias (2009) and limited kernel flexibility Wilson et al. (2016); Liu et al. (2021). In contrast, Neural Networks (NNs) Schmidhuber (2015); Nielsen (2015) are highly flexible and scalable but lack inherent mechanisms for uncertainty modeling Blundell et al. (2015); Pearce et al. (2020); Gawlikowski et al. (2023). NPs address these limitations by modeling distributions over functions using a neural network-based framework. They approximate a stochastic process $F:X\to Y$ through finite-dimensional marginals, parameterized by a latent variable $z$ to capture global uncertainty. Given context observations $(x_{\mathbb{C}},y_{\mathbb{C}})$ and target inputs $x_{\mathbb{T}}$, NPs generate predictive distributions over $y_{\mathbb{T}}$ via a conditional latent model.
$p(y_{\mathbb{T}},z|x_{\mathbb{T}},x_{\mathbb{C}},y_{\mathbb{C}})=p(z|x_{\mathbb{C}},y_{\mathbb{C}})\prod_{i=1}^{|\mathbb{T}|}p(y_{\mathbb{T},i}|x_{\mathbb{T},i},z)$ (62)
@@
3. Decoder $g$: Predicts the target outputs $y_{\mathbb{T}}=g(x_{\mathbb{T}},z)$, conditioned on the latent variable $z$ and the target inputs $x_{\mathbb{T}}$.
@@
$\log p(y_{\mathbb{T}}|x_{\mathbb{C}},y_{\mathbb{C}},x_{\mathbb{T}})\geq\mathbb{E}_{q(z|x_{\mathbb{C}},y_{\mathbb{C}})}\left[\sum_{i\in\mathbb{T}}\log p(y_{\mathbb{T},i}|z,x_{\mathbb{T},i})+\log\frac{p(z|x_{\mathbb{C}},y_{\mathbb{C}})}{q(z|x_{\mathbb{C}},y_{\mathbb{C}})}\right]$ (63)
where $q(z|x_{\mathbb{C}},y_{\mathbb{C}})$ is the variational posterior distribution parameterized by a neural network, and $p(z|x_{\mathbb{C}},y_{\mathbb{C}})$ is the conditional prior.
Appendix D NDP Review
Formally, given a function $f:\mathbb{R}^{D}\to\mathbb{R}$, an NDP learns a generative distribution over observed data pairs $(x,y)$, where inputs $x\in\mathbb{R}^{N\times D}$ and outputs $y=f(x)\in\mathbb{R}^{N}$. Unlike standard NPs, NDPs Dutordoir et al. (2023) do not explicitly require a partitioning into context and target sets during training; all points are jointly modeled. In supervised learning setting, the NDP modeling framework consists of two stochastic processes:
@@ Appendix D NDP Review
NDPs employ a denoising score matching objective  Hyvärinen & Dayan (2005); Song et al. (2021); Huang et al. (2021), training the noise model $\epsilon_{\theta}$ by minimizing the discrepancy between predicted noise and actual noise $\epsilon\sim\mathcal{N}(0,I)$:
@@
At test time, NDPs draw samples from the conditional distribution $p(y_{\mathbb{T},0}\mid x_{\mathbb{T}},D)$, where $D=(x_{\mathbb{C}}\in\mathbb{R}^{M\times D},y_{\mathbb{C},0}\in\mathbb{R}^{M})$ is the context observations.
The conditional sampling proceeds as follows. First, sample the initial target noise: $y_{\mathbb{T},T}\sim\mathcal{N}(0,I).$ For each diffusion timestep $t=T,\ldots,1$, proceed with:
@@
    $y_{\mathbb{C},t}\sim\mathcal{N}\left(\sqrt{\bar{\alpha}_{t}}y_{\mathbb{C},0},(1-\bar{\alpha}_{t})I\right),$ (69)
@@
    $y_{t}=\{y_{\mathbb{T},t},y_{\mathbb{C},t}\},\quad x=\{x_{\mathbb{T}},x_{\mathbb{C}}\}.$ (70)
@@
Appendix E Noise Model Architecture
To ensure that our model remains consistent with the structural properties of stochastic processes and to guarantee fair experimental comparisons, we adopt the same noise model architecture as NDPs, namely the Bi-Dimensional Attention Block Dutordoir et al. (2023).
As decsribed in Figure 1, this architecture is designed to encode two key symmetries:
@@ Appendix E Noise Model Architecture
Proposition 1 (Equivariance (Dutordoir et al., 2023, Prop. 4.1))
@@
Proposition 2 ( Equivariance (Dutordoir et al., 2023, Prop. 4.2))
@@
By directly encoding these properties into the noise model architecture, NDPs and NBPs ensure that the predicted outputs $\{y_{t}^{1},\dots,y_{t}^{N}\}$ at each timestep $t$ form an exchangeable set of random variables, consistent with the Kolmogorov Extension Theorem (KET). This is critical for defining a valid stochastic process over functions.
Appendix F More Details for Experiments
F.1 Baseline Implementation and Evaluation Metrics
To provide a comprehensive comparison, we implement NPs Garnelo et al. (2018b) , ANPs Kim et al. (2019), and ConvNPs Gordon et al. (2019) using the official NP-Family repository Dubois et al. (2020), with all hyperparameters set to the recommended default values. For NDP Dutordoir et al. (2023), we directly utilize the official repositories. Meanwhile, to ensure a fair comparison, our NBP model adopts the same Bi-Dimensional Attention Block architecture and hyperparameters as the baseline NDP. The implementation details are provided in the supplementary materials. All models are retrained on the experimental datasets to ensure consistent metric evaluation and visualization. We evaluate model performance using two primary metrics: Mean Squared Error (MSE) and Negative Log-Likelihood (NLL). All experiments are conducted on a single NVIDIA RTX 4090 GPU.
@@ F.1 Baseline Implementation and Evaluation Metrics
F.2 Details of EEG Dataset Regression Task
@@
For our study, we focus on signals from 7 frontal electrodes: FZ, F1, F2, F3, F4, F5, and F6. This selection yields a total of 7,632 multivariate time series, each comprising 256 time steps across 7 channels. These signals exhibit strong temporal dynamics and inter-channel correlations, making the dataset well-suited for evaluating the generalization and modeling capabilities of multi-output meta-learning models. The data is publicly available from the UCI Machine Learning Repository, with collection details described in Zhang et al. (1995). Figure 6 illustrates the signals recorded from these seven channels for a single trial of one subject.
@@
Within each trial, we randomly select 3 out of the 7 channels and mask partial segments of these channels to simulate missing data. This setup supports evaluation across the following three tasks:
Interpolation: Recovering locally missing values within the observed time range.
Reconstruction: Reconstructing entire masked regions of a target channel using the remaining channels as context.
Forecasting: Extrapolating future signal trajectories based on current observations.
Each input is represented as an index vector $\mathbf{x}_{e}=(i_{t},i_{c})$, where $i_{t}$ denotes the time step and $i_{c}$ the channel index. The corresponding output is the voltage signal $\mathbf{y}_{e}$.
All models were trained for 1,000 iterations on the training set. Evaluation metrics include Mean Squared Error (MSE) and Negative Log-Likelihood (NLL). The Neural Bridge Process (NBP) employs a 5-layer Bi-Dimensional Attention Block with hidden dimension 64 and 8 attention heads. The model was trained using a fixed random seed of 42 to ensure reproducibility.
For models incorporating diffusion-based generation (such as DNP and NBP), we adopt a cosine noise schedule with the following parameters: $\beta_{\text{start}}=0.0003$, $\beta_{\text{end}}=0.5$, and 500 diffusion timesteps. These settings are applied consistently across the forward and reverse processes in all diffusion-based models. Additional training hyperparameters are as follows:
  • For NBP and DNP base, the learning rate was set to $2\cdot 10^{-5}$;
@@
The evaluation results, summarized in Table 1 of the main text, demonstrate that NBP consistently outperforms existing baseline models across all three tasks, highlighting its superior modeling capacity for highly correlated multichannel temporal data.
F.3 Details of Image Regression Task
@@
Our experimental protocol—including the denoising network architecture, training schedule, optimizer configuration, and random seed—closely follows the setup used for Neural Diffusion Processes (NDPs), ensuring a fair and consistent comparison. The core architecture of the NBP model consists of 7 layers, each with hidden dimension 64 and 8 attention heads. Sparse attention is not used in these experiments.
@@
The model is trained for 10 epochs using a batch size of 32. The optimizer follows a warmup and decay schedule:
@@
  • Warmup: 20 epochs, decay over 200 epochs
@@
Each prediction is averaged over 9 conditional samples during testing. The evaluation batch size is set to 9, with 128 samples drawn per image for final averaging. All pixel values are normalized to the $[0,1]$ range. The model was trained and evaluated with a fixed random seed of 42.

### v2

@@ 1. The ELBO decomposes into:
2. Sample from $q(y_{t}|y_{0},x)$: Using the forward marginal in Proposition 4.1, we can write:
@@ 5. Final Noise-Prediction Loss: The KL terms simplify to a weighted $L_{2}$ loss on the noise:
This proves Eq. (14), and the denoising network $\epsilon_{\theta}(y_{t},t,x)$ is self-consistent with respect to the condition on $x$.
B.4 Proof of Theorem 4.4
For fixed $y_{0}$, the two NBP marginals have the same covariance $(1-\bar{\alpha}_{t})I$ and means
$m_{t}(x)=\sqrt{\bar{\alpha}_{t}}y_{0}+\bar{\gamma}_{t}x,\qquad m_{t}(x^{\prime})=\sqrt{\bar{\alpha}_{t}}y_{0}+\bar{\gamma}_{t}x^{\prime}.$
For two Gaussians with identical covariance, the squared 2-Wasserstein distance equals the squared Euclidean distance between their means. Therefore,
$W_{2}^{2}=\|m_{t}(x)-m_{t}(x^{\prime})\|_{2}^{2}=\bar{\gamma}_{t}^{2}\|x-x^{\prime}\|_{2}^{2},$
which proves the claim. For NDP, the forward marginal does not contain $x$, so the two marginals are identical and the distance is zero.
B.5 Proof of Theorem 4.5
Conditioning on $y_{0}$, the NBP forward marginal can be written as a Gaussian channel
$y_{t}-\sqrt{\bar{\alpha}_{t}}y_{0}=\bar{\gamma}_{t}x+\eta_{t},\qquad\eta_{t}\sim\mathcal{N}(0,(1-\bar{\alpha}_{t})I),$
where $\eta_{t}$ is independent of $x$. If $x\sim\mathcal{N}(0,\Sigma_{x})$, the mutual information of this linear Gaussian channel is
$I(x;y_{t}\mid y_{0})=\frac{1}{2}\log\det\left(I+(1-\bar{\alpha}_{t})^{-1}\bar{\gamma}_{t}^{2}\Sigma_{x}\right).$
For NDP, $y_{t}=\sqrt{\bar{\alpha}_{t}}y_{0}+\eta_{t}$ conditional on $y_{0}$, so $y_{t}$ is independent of $x$ and the mutual information is zero.
B.6 Proof of Proposition 4.6
The NBP noised state is
$y_{t}(x)=\sqrt{\bar{\alpha}_{t}}y_{0}+\bar{\gamma}_{t}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon.$
Applying the chain rule to $\ell_{t}(\theta,x)=\|\epsilon_{\theta}(y_{t}(x),x,t)-\epsilon\|_{2}^{2}$ gives
$\frac{d\ell_{t}}{dx}=\frac{\partial\ell_{t}}{\partial x}+\frac{\partial\ell_{t}}{\partial y_{t}}\frac{\partial y_{t}}{\partial x}=\frac{\partial\ell_{t}}{\partial x}+\bar{\gamma}_{t}\frac{\partial\ell_{t}}{\partial y_{t}}.$
The first term is the ordinary conditioning path through the denoiser input $x$, while the second term is induced by the input-anchored forward state. In NDP, the forward process does not contain $x$, so $\partial y_{t}/\partial x=0$ and the second term disappears.
@@ Appendix C Background: Neural Processes
Neural Processes (NPs) [14, 13, 23, 28] combine the expressiveness of neural networks with the probabilistic reasoning of Gaussian Processes (GPs) [37]. While GPs offer principled uncertainty quantification, they suffer from poor scalability [42, 45] and limited kernel flexibility [46, 27]. In contrast, Neural Networks (NNs) [39, 33] are highly flexible and scalable but lack inherent mechanisms for uncertainty modeling [1, 35, 15]. NPs address these limitations by modeling distributions over functions using a neural network-based framework. They approximate a stochastic process $F:X\to Y$ through finite-dimensional marginals, parameterized by a latent variable $z$ to capture global uncertainty. Given context observations $(x_{\mathcal{C}},y_{\mathcal{C}})$ and target inputs $x_{\mathcal{T}}$, NPs generate predictive distributions over $y_{\mathcal{T}}$ via a conditional latent model.
$p(y_{\mathcal{T}},z|x_{\mathcal{T}},x_{\mathcal{C}},y_{\mathcal{C}})=p(z|x_{\mathcal{C}},y_{\mathcal{C}})\prod_{i=1}^{|\mathcal{T}|}p(y_{\mathcal{T},i}|x_{\mathcal{T},i},z)$ (60)
@@
3. Decoder $g$: Predicts the target outputs $y_{\mathcal{T}}=g(x_{\mathcal{T}},z)$, conditioned on the latent variable $z$ and the target inputs $x_{\mathcal{T}}$.
@@
$\log p(y_{\mathcal{T}}|x_{\mathcal{C}},y_{\mathcal{C}},x_{\mathcal{T}})\geq\mathbb{E}_{q(z|x_{\mathcal{C}},y_{\mathcal{C}},x_{\mathcal{T}},y_{\mathcal{T}})}\big[\sum_{i\in\mathcal{T}}\log p(y_{\mathcal{T},i}|z,x_{\mathcal{T},i})+\log\frac{p(z|x_{\mathcal{C}},y_{\mathcal{C}})}{q(z|x_{\mathcal{C}},y_{\mathcal{C}},x_{\mathcal{T}},y_{\mathcal{T}}))}\big]$ (61)
where $q(z|x_{\mathcal{C}},y_{\mathcal{C}},x_{\mathcal{T}},y_{\mathcal{T}})$ is the variational posterior distribution parameterized by a neural network, and $p(z|x_{\mathcal{C}},y_{\mathcal{C}})$ is the conditional prior.
Appendix D Input-output Dimensional Alignment
Real-world datasets often have mismatched input and output dimensions. Figure 3 summarizes how NBP handles this case. To apply the bridge in the output space, we introduce an anchor map
$a_{\psi}(x)=P_{\psi}(x),\qquad P_{\psi}:\mathbb{R}^{D_{x}}\rightarrow\mathbb{R}^{D_{y}},$ (62)
which maps each input location to an output-space anchor. The bridge term is then formed with $a_{\psi}(x)$ rather than the raw input $x$:
$q(y_{t}\mid y_{t-1},x)=\mathcal{N}\left(y_{t};\sqrt{1-\beta_{t}}y_{t-1}+\gamma_{t}a_{\psi}(x),\beta_{t}I\right).$ (63)
This design separates two roles of the input. The raw input $x$ is still passed to the denoising network as conditioning information, while $a_{\psi}(x)$ is used only to define an output-space bridge anchor. In image regression, for instance, $x$ is a two-dimensional normalized pixel coordinate and $y$ is a three-dimensional RGB value. We therefore parameterize $P_{\psi}$ as a lightweight coordinate-to-output anchor network: the coordinates are first positionally encoded, then passed through an MLP anchor head that outputs a vector in the RGB space. The anchor head is trained jointly with the denoising model using the main denoising objective and an auxiliary anchor reconstruction loss $\|a_{\psi}(x)-y_{0}\|^{2}$. This auxiliary term is used only to make the bridge anchor live in the correct output space; it does not change the denoising architecture, the number of reverse sampling steps, or the number of training updates. When $D_{x}=D_{y}$, as in the 1D GP experiments, $P_{\psi}$ can reduce to the identity map.
[…]
Figure 3: Learned input-output alignment for NBP when $D_{x}\neq D_{y}$. The raw input remains denoiser conditioning, while $P_{\psi}$ maps it into the output space before it enters the bridge trajectory. In image regression, this maps 2D pixel coordinates to RGB-valued anchors.
Appendix E NDP Review
Formally, given a function $f:\mathbb{R}^{D}\to\mathbb{R}$, an NDP learns a generative distribution over observed data pairs $(x,y)$, where inputs $x\in\mathbb{R}^{N\times D}$ and outputs $y=f(x)\in\mathbb{R}^{N}$. Unlike standard NPs, NDPs [10] do not explicitly require a partitioning into context and target sets during training; all points are jointly modeled. In supervised learning setting, the NDP modeling framework consists of two stochastic processes:
@@ Appendix E NDP Review
NDPs employ a denoising score matching objective  [22, 43, 21], training the noise model $\epsilon_{\theta}$ by minimizing the discrepancy between predicted noise and actual noise $\epsilon\sim\mathcal{N}(0,I)$:
@@
At test time, NDPs draw samples from the conditional distribution $p(y_{\mathcal{T},0}\mid x_{\mathcal{T}},D)$, where $D=(x_{\mathcal{C}}\in\mathbb{R}^{M\times D},y_{\mathcal{C},0}\in\mathbb{R}^{M})$ is the context observations.
The conditional sampling proceeds as follows. First, sample the initial target noise: $y_{\mathcal{T},T}\sim\mathcal{N}(0,I).$ For each diffusion timestep $t=T,\ldots,1$, proceed with:
@@
    $y_{\mathcal{C},t}\sim\mathcal{N}\left(\sqrt{\bar{\alpha}_{t}}y_{\mathcal{C},0},(1-\bar{\alpha}_{t})I\right),$ (69)
@@
    $y_{t}=\{y_{\mathcal{T},t},y_{\mathcal{C},t}\},\quad x=\{x_{\mathcal{T}},x_{\mathcal{C}}\}.$ (70)
@@
Appendix F Noise Model Architecture
To ensure that our model remains consistent with the structural properties of stochastic processes and to guarantee fair experimental comparisons, we adopt the same noise model architecture as NDPs, namely the Bi-Dimensional Attention Block [10].
Figure 4: Noise model architecture employed at each step of Neural Bridge Processes. The greyed-out box highlights the bi-dimensional attention block inherited from NDP.
As described in Figure 4, this architecture is designed to encode two key symmetries:
@@ Appendix F Noise Model Architecture
Proposition F.1 (Equivariance [Prop. 4.1][10]).
@@
Proposition F.2 ( Equivariance [Prop. 4.2][10]).
@@
By encoding these structural properties, NDP and NBP generate, at each step, a set of exchangeable random variables— a structural requirement commonly used in NDP models to construct consistent finite-dimensional conditional distributions. However, as with NDP, our method does not attempt to satisfy the strict marginal consistency conditions required by the Kolmogorov Extension Theorem, even though the structure allows the model to empirically approximate consistency during meta-learning.
Appendix G More Details for Experiments
G.1 Baseline Implementation and Evaluation Metrics
To provide a comprehensive comparison, we implement NPs [14] , ANPs [23], and ConvNPs [16] using the official NP-Family repository [9], with all hyperparameters set to the recommended default values. For NDP [10], we directly utilize the official repositories. Meanwhile, to ensure a fair comparison, our NBP model adopts the same Bi-Dimensional Attention Block architecture and hyperparameters as the baseline NDP. The implementation details are provided in the supplementary materials. All models are retrained on the experimental datasets and evaluated with the same seed set to ensure consistent metric evaluation and visualization. We evaluate model performance using two primary metrics: Mean Squared Error (MSE) and Negative Log-Likelihood (NLL). Experiments are conducted on NVIDIA RTX A6000 GPUs with 48GB memory; each reported run uses a single GPU unless otherwise stated.
@@ G.1 Baseline Implementation and Evaluation Metrics
G.2 Details of EEG Dataset Regression Task
@@
For our study, we focus on signals from 7 frontal electrodes: FZ, F1, F2, F3, F4, F5, and F6. This selection yields a total of 7,632 multivariate time series, each comprising 256 time steps across 7 channels. These signals exhibit strong temporal dynamics and inter-channel correlations, making the dataset well-suited for evaluating the generalization and modeling capabilities of multi-output meta-learning models. The data is publicly available from the UCI Machine Learning Repository, with collection details described in [51]. Figure 5 illustrates the signals recorded from these seven channels for a single trial of one subject.
@@
Within each trial, we randomly select 3 out of the 7 channels and mask partial segments of these channels to simulate missing data. This setup supports three complementary prediction regimes. Interpolation evaluates whether the model can recover locally missing values within the observed time range. Reconstruction asks the model to infer masked regions of a target channel using the remaining channels as context, thereby testing cross-channel dependency modeling. Forecasting evaluates extrapolation to future signal trajectories based on currently observed measurements. Each input is represented as an index vector $\mathbf{x}_{e}=(i_{t},i_{c})$, where $i_{t}$ denotes the time step and $i_{c}$ the channel index. The corresponding output is the voltage signal $\mathbf{y}_{e}$.
All models were trained for 1,000 iterations on the training set. Evaluation metrics include Mean Squared Error (MSE) and Negative Log-Likelihood (NLL). The Neural Bridge Process (NBP) employs a 5-layer Bi-Dimensional Attention Block with hidden dimension 64 and 8 attention heads. We train and evaluate each model over multiple random seeds and report the mean performance, using the same seed set for all compared methods.
For models incorporating diffusion-based generation (such as NDP and NBP), we adopt a cosine noise schedule with the following parameters: $\beta_{\text{start}}=0.0003$, $\beta_{\text{end}}=0.5$, and 500 diffusion timesteps. These settings are applied consistently across the forward and reverse processes in all diffusion-based models. Additional training hyperparameters are as follows:
  • For NBP and NDP base, the learning rate was set to $2\cdot 10^{-5}$;
@@
The evaluation results, summarized in Table 2 of the main text, show that NBP obtains lower NLL and MSE than the compared baselines across the three EEG prediction tasks.
Table 3: Calibration summary on the EEG interpolation setting. ECE and sharpness are estimated from predictive samples; lower ECE indicates better calibration, lower sharpness indicates more concentrated predictive uncertainty, and 90% coverage is best when close to 0.90.
Model NLL ($\downarrow$) MSE($\times 10^{-2}$) ($\downarrow$) ECE ($\downarrow$) Sharpness $\bar{\sigma}$ ($\downarrow$) 90% Coverage
NP 1.66 0.52 0.182 0.91 0.71
ANP 0.47 0.25 0.134 0.67 0.79
ConvNP 0.44 0.40 0.128 0.63 0.80
NDP -2.46 0.18 0.089 0.38 0.83
SNP -3.19 0.16 0.061 0.29 0.88
GEOMNDP -2.48 0.18 0.087 0.37 0.83
NBP (ours) -3.55 0.15 0.047 0.24 0.91
G.3 Details of Image Regression Task
(a) Ground Truth
(b) Corrupted Context
(c) Model Prediction
Figure 6: Qualitative results on the CelebA $\times64$ image regression task.
@@ G.3 Details of Image Regression Task
Our experimental protocol—including the denoising network architecture, training schedule, optimizer configuration, and seed set—closely follows the setup used for Neural Diffusion Processes (NDPs), ensuring a fair and consistent comparison. The core architecture of the NBP model consists of 7 layers, each with hidden dimension 64 and 8 attention heads. Sparse attention is not used in these experiments.
@@
The model is trained using a batch size of 32. The optimizer follows the same learning-rate schedule as the NDP baseline:
@@
  • Warmup and cosine decay are applied following the NDP implementation.
@@
Each prediction is averaged over 9 conditional samples during testing. The evaluation batch size is set to 9, with 128 samples drawn per image for final averaging. All pixel values are normalized to the $[0,1]$ range. Results are averaged over multiple random seeds using the same seed set across compared methods.